Find three cnsecutive odd integers such that twice the first, decreased by the second is 35. : Find three cnsecutive odd integers such that twice the first, decreased by the second is 35.
Answer by nerdybill(1042)

Find three cnsecutive odd integers such that twice the first, decreased by the second is 35.
.
Let x = first consecutive odd integer
x + 2 = second consecutive odd integer
x + 4 = third consecutive odd integer
.
2x - (x + 2) = 35
2x - x - 2 = 35
x - 2 = 35
x = 37
.
Therefore, the three consecutive odd integers are:
37, 39 and 41